OpenStudy (anonymous):

Suppose a triangle has sides a, b, and c, and that a2 + b2 < c2. Let be the measure of the angle opposite the side of length c. Which of the following must be true? Check all that apply.

OpenStudy (anonymous):

\[A. cos \theta\] <0 B. the triangle is a right triangle C. the triangle is not a right triangle \[D. \theta \] is an obtuse angle help me please

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

I need more help I do not understand

OpenStudy (raden):

use the cosine rule

OpenStudy (raden):

cos C = (a^2 + b^2 - c^2)/(2ab) we knowed that a^2 + b^2 < c^2 obviously, the numerator be negative and the denominator is positive. cos C = -/+ cos C = - so, <C is in the 2nd quadrant (more than 90 degrees) finally, it is an obtuse triangle

OpenStudy (anonymous):

so the answer is only D

OpenStudy (raden):

wait, is the theta the measure of the angle opposite the side of length c ?

OpenStudy (raden):

check the original question.... let .?.. (is it theta)

OpenStudy (anonymous):

you see the picture

OpenStudy (raden):

aha, i thought that is the theta :) you just miss it after you type let ........

OpenStudy (raden):

cos C = cos θ like i said above : cos C = - it means that cos θ < 0 (A is correct, and D is correct too)

OpenStudy (anonymous):

Suppose a triangle has sides a, b, and c, and that a2 + b2 < c2. Let be the measure of the angle opposite the side of length c. Which of the following must be true? Check all that apply. A. The triangle is not a right triangle. B. The triangle is a right triangle. C. cos < 0 D. is an acute angle.